OpenStudy (anonymous):

Two sound waves ofthe same frequency having amplitudes of 3 cm and I cm respectively are moving towards each other in the same straight line, The resultant wave produced will vary in loudness. The ratio of the maximum to minimum intensity is (A) 4/1 (B) 9/2 (C) 6/1 (D) 9/1

OpenStudy (anonymous):

What don't you understand?

OpenStudy (anonymous):

Intensity is directly proportional to the square of amplitude. I=(Amplitude)^2 so the ration I(max):I(min) =3^2/1^2=9:1

OpenStudy (anonymous):

*so the ratio I(max):I(min) =3^2/1^2=9:1

OpenStudy (anonymous):

interference would cause the max amplitude to be (3+1)=4 and min amplitude to be (3-1)=2 intensities would then be: max (4)^2, min (2)^2. Correct?

OpenStudy (anonymous):

The question has not mentioned about interference occurring here. Besides your answer is not in the option as well. In case they interfere, we further need to know the phase difference between the waves to determine the maximum and minimum intensity of the resulting wave.

OpenStudy (anonymous):

1)They are traveling toward each other, so will interfere. 2) The answer 16/4=4/1, is listed here. 3) The waves are of the same frequency, and so will have points of maximum and minimum interference. What other information is necessary?

OpenStudy (anonymous):

@ nsutton.. Yes you are correct. Sorry! I misunderstood the question..

OpenStudy (anonymous):

All good. Happens to me all the time. :)